OpenStudy (anonymous):

(x-4)/(sqroot(x)-2) multiplied by its reciprocal

OpenStudy (anonymous):

any expression multiplied by its reciprocal = 1 eg reciprocal of x = 1/x x * 1/x = 1

OpenStudy (campbell_st):

if the question is \[(x - 4)/(\sqrt{x} -2) \times (\sqrt{x} -2)/(x-4) \] the answer is 1.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think i used the wrong term... i have to multiply but the opposite of the bottom

OpenStudy (anonymous):

but=by

OpenStudy (anonymous):

the opposite of the bottom?

OpenStudy (anonymous):

ya so its (x-4)/(sqroot(x)-2)*(sqroot(x)+2)/(sqroot(x)+2)

OpenStudy (anonymous):

hmm maybe you want to simplify by rationalising the denominator i e multoplying top and bottom by the conjugate of sqrtx - 2 which is sqrtx + 2 so its (x-4)*(sqrtx + 2) (x-4)(sqrtx+2) -------------- = ------------ = sqrtx + 2 (sqrtx-2)(sqrtx+2) ( x - 4)

OpenStudy (anonymous):

yes conjugate is the word i was looking for

OpenStudy (anonymous):

the (x-4)'s cancel out and you're left with sqrtx + 2

OpenStudy (anonymous):

glad to be of help

OpenStudy (anonymous):

thank you